OpenStudy (anonymous):

Y~B(16,0.8) P(Y>15.36) Anyone???

OpenStudy (mathmate):

Normal distribution?

OpenStudy (anonymous):

yup

OpenStudy (mathmate):

Do you have a table of distributions of z vs probability? z = (15.36-16)/0.8 = - 0.8

OpenStudy (anonymous):

Yup, I do, but it doesn't go to decimal point. So, if I were to look for n=16, there wouldn't be any for p=15.36

OpenStudy (anonymous):

The answer is 0.1407, not sure at all how to get that if the table doesn't have it...

OpenStudy (mathmate):

You'd want to convert the number in terms of mean and sd, kind of to standardize it. The mean is 16, the value is 15.36, so the number of sd from the mean is then z=(16-15.36) divided by the sd = 0.64/0.8 = 0.8 The tables give probabilities in terms of the z value, which is 0.7881 corresponding to z=0.8

OpenStudy (anonymous):

Yes, it does give it to 0.8 as P. Sorry, it's the X value that isn't shown. 15.36 is the x value, n=16, P=0.8. I have 16 and 0.8 but not 15.36... It doesn't go in decimals... And the answer is 0.1407... again, not sure how to solve it...

OpenStudy (mathmate):

Is 16 the mean, and 0.8 the sd?

OpenStudy (anonymous):

I mean 0.8 is the probability

OpenStudy (mathmate):

Ah.. let me see..

OpenStudy (anonymous):

15.36 is the mean + the variation

OpenStudy (mathmate):

There's something I do not understand, so the mean is 16. 15.36 is an observation, say x. and 0.8 is a probability... of what?

OpenStudy (anonymous):

This is the full question: For the random variable Y, for which Y ~ B(16,0.8), calculate P(Y>mean+variation)

OpenStudy (anonymous):

So, 0.8 is just the probability of something occuring... (it could be anything)

OpenStudy (mathmate):

So the variable does not have to be normally distributed?

OpenStudy (anonymous):

It's binomial distribution... Which is the same as normal distribution, isn't it? I'm just doing statistics 1, so it should be normally distributed

OpenStudy (anonymous):

Sorry, internet connection problems.

OpenStudy (mathmate):

No, binomial distribution has a different set of values of probabilities.

OpenStudy (anonymous):

Ah Ok. So, How would you do that here?

OpenStudy (mathmate):

So 16 is the number of trials, and 15.36 is the minimum success required, and 0.8 is the probability of success, right?

OpenStudy (anonymous):

Yup. That's correct.

OpenStudy (anonymous):

Sorry, electricity went.

OpenStudy (mathmate):

Have you used the formula to calculate the probability, instead of looking up tables?

OpenStudy (mathmate):

If n=no. of trials k=min. success required, and p=probability of success, then P(k,n,p)=(n,k)(p^k)(1-p)^k Does that sound familiar to you?

OpenStudy (anonymous):

No, it isn't familiar to me, but it does make a lot more sense. Do I just plug in the values?

OpenStudy (mathmate):

I am not too familiar with this distribution, but try it anyway.

OpenStudy (anonymous):

OK. I'll tell you if it works

OpenStudy (mathmate):

Thks.

OpenStudy (anonymous):

but... how do I do (n,k)?

OpenStudy (mathmate):

That's probably where they require an integer. (n,k) is n!/((n!)(n-k)!) where n! means the factorial of n.

OpenStudy (anonymous):

Ah Ok.

OpenStudy (mathmate):

I made a mistake with the formula I gave, it should read P(k,n,p)=(n,k)(p^k)(1-p)^(n-k) I plugged in the formula with k=15.36 and got 0.0766.

OpenStudy (mathmate):

If you are in a rush, you may want to submit a new thread and specify explicitly it's a binomial distribution. I should have caught it when you put B(16,0.8), but didn't.

OpenStudy (anonymous):

Ah Ok. Yes, it was wrong when I did the first equation... I'm not in a rush. It's just something I have to understand. It's my internet which keeps falling away that's the problem.

OpenStudy (anonymous):

Yes, it's not that answer, but I give you a medal for trying :)

OpenStudy (mathmate):

Sorry, I wish I were more familiar with cumulative binomial distributions. If I found anything later on, I'll post it here.

OpenStudy (anonymous):

I'll repost... but last time I did that, no one went to answer it... I guess Binomial Distribution scares people, and they'd rather go for the more simpler Distribution or equations... so now I tend not to write it as Binomial distribution.. but I guess I will :)

OpenStudy (anonymous):

Are you good at normal probability distributions?